OpenStudy (anonymous):

for the given polynomial P(x) and the given c, use the remainder theorem to fid P(c). P(x)=3x^3-7x^2-2x+5 ; c=-3

OpenStudy (experimentx):

-133

OpenStudy (anonymous):

how did you get that?

OpenStudy (experimentx):

put the value x=-3, it seems that x+3 is not factor of P(x)

OpenStudy (anonymous):

so for this one P(x)=4x^4+x^2-2 ; c= -1 if it doesnt equal 0 its not a factor?

OpenStudy (anonymous):

or does it have to equal the the value of c?

OpenStudy (experimentx):

yup ...

OpenStudy (experimentx):

no ... it must be zero if x+1 is a factor of P(x)

OpenStudy (experimentx):

let me say again ... if x+a is a factor of P(x) then p(-a) must be equal to zero

OpenStudy (anonymous):

ohh ok when plugging in c to p(x) it must equal 0 for it to be a factor

OpenStudy (experimentx):

for example we have P(x) = x^2 - 3x + 2 now we have P(-1) = 0, so, x+1 is factor of P(x) and so is x-2, but P(5) != 0 therefore x-5 is not factor of P(x)

OpenStudy (anonymous):

ohh can you do just this one P(x)=4x^4+x^2-2 ; c= -1

OpenStudy (experimentx):

P(-1) = 4 + 1 - 2 = 3 != 0 so x+1 is not a factor of P(x)

OpenStudy (anonymous):

i get it! Thank you